OpenStudy (anonymous):

help!! What translation has taken place in the graph of f(x)=x^3-4from the graph of the parent function? A.f(x) was shifted up four units B.f(x) was shifted down four units C.f(x) was shifted right four units D.f(x) was shifted left four units

OpenStudy (jdoe0001):

A(Bx + C) + D C = horizontal translation, C > 0, to the left, C < 0,, to the right D = vertical translation, D > 0, up, D < 0, down

OpenStudy (anonymous):

so what would that make it?

OpenStudy (jdoe0001):

well, lemme rerwrite yours \(\large \bf f(x)=x^3-4 \implies f(x) = 1(1x^3+ 0) + 4\) so, what do you think?

OpenStudy (anonymous):

it shifted left four units

OpenStudy (jdoe0001):

lemme rewrite it again $$\bf f(x)=x^3-4 \\ f(x) = 1(1x^3+ 0) + 4 \\ A(Bx + C) + D$$

OpenStudy (anonymous):

right..

OpenStudy (jdoe0001):

so, is not a horizontal shift

OpenStudy (anonymous):

so it's A

OpenStudy (jdoe0001):

anyhow, is -4 what you have, yes, is vertical shift, I by mistake put a +4 but you have $$ \bf f(x)=x^3-4 \\ f(x) = 1(1x^3+ 0) - 4 \\ A(Bx + C) + D $$ so it's vertically shifted DOWN by 4 units